OpenStudy (yacoub1993):

Solve the absolute value equation or indicate that the equation has no solution. │x + 6│ = 9 A. {3} B. {-15, 3} C. {-3, 15} D. ø

hartnn (hartnn):

when |a| =b either a=b or a=-b so what do u get for |x+6| = 9 ?

OpenStudy (yacoub1993):

it is B WRITE

hartnn (hartnn):

you get x+6 = 9 or x+6=-9 can you solve these separately ?

hartnn (hartnn):

and yes, its B :)

OpenStudy (yacoub1993):

X+6=9 X=9-6 X=3 X+6=-9 X=-9-6 X=-15

OpenStudy (yacoub1993):

RIGHT

hartnn (hartnn):

yes, thats correct solution , good!

OpenStudy (yacoub1993):

thnxs... i will post another question

hartnn (hartnn):

welcome ^_^